How far away from each other would two metal plates, 2 square meters in area each, must be in order to create a capacitance of 1 μF? Assume that the plates are separated by air.


1
Expert's answer
2022-03-15T10:34:56-0400

The capacitance of parallel plate capacitor

"C=\frac{\epsilon_0 A}{d}"

Hence, the separation between plates

"d=\frac{\epsilon_0 A}{C}=\frac{8.85*10^{-12}*2.00}{1.00*10^{-6}}=1.77*10^{-5}\:\rm m"